# Statistical Inference for Explainable Boosting Machines

### v1

@@ Abstract
Explainable boosting machines (EBMs) are popular “glass-box” models that learn a set of univariate functions using boosting trees. These achieve explainability through visualizations of each feature’s effect. However, unlike linear model coefficients, uncertainty quantification for the learned univariate functions requires computationally intensive bootstrapping, making it hard to know which features truly matter. We provide an alternative using recent advances in statistical inference for gradient boosting, deriving methods for statistical inference as well as end-to-end theoretical guarantees. Using a moving average instead of a sum of trees (Boulevard regularization) allows the boosting process to converge to a feature-wise kernel ridge regression. This produces asymptotically normal predictions that achieve the minimax-optimal MSE for fitting Lipschitz GAMs with $p$ features of $O(p\,n^{-2/3})$, successfully avoiding the curse of dimensionality. We then construct prediction intervals for the response and confidence intervals for each learned univariate function with a runtime independent of the number of datapoints, enabling further explainability within EBMs.
@@ 1 Introduction
Modern machine learning methods have proven to be highly successful, achieving impressively high performance even for complicated prediction tasks. Still, this comes at a cost. These methods have been critiqued for their lack of transparency, often described as “black-box” models. As such, much work has been done in an attempt to provide explanations for their behavior. [^1] Popular methods include LIME (Ribeiro et al., 2016), SHAP (Lundberg and Lee, 2017), ALE (Apley and Zhu, 2020) and partial dependence (Friedman, 2001) plots.Another line of work attempts to convince practitioners to instead use inherently transparent models that are algebraically simple enough to be inspected by humans (Rudin, 2019). Neither is optimal. Explaining black-box models is a difficult problem (Hooker et al., 2021; Fryer et al., 2021), and inherently interpretable models can sacrifice performance.
@@
In light of this, Explainable Boosting Machines (EBM) (Lou et al., 2012; Nori et al., 2019) have emerged as a popular “glass-box” model to strike a balance between both. An EBM learns a Generalized Additive Model (GAM) (Hastie and Tibshirani, 1986): $\mathbf{f}(\mathbf{x})=\sum_{k=1}^{p}\mathbf{f}^{(k)}(\mathbf{x}^{(k)})$, where each $\mathbf{f}^{(k)}$ is assumed to be a univariate function of the $k$-th feature. However, instead of fitting splines, it instead fits a set of univariate gradient boosting regression trees. The tree-based boosting procedure attempts to recover the expressiveness of state-of-the-art algorithms (Chen and Guestrin, 2016; Ke et al., 2017), while the decomposition into additive components allows the measurement of the marginal contribution of each feature like linear models. While EBMs impose strong structural constraints on the resulting prediction, their ability to reproduce nonlinearities, along with appropriate choices of additive structure, provides comparable performance to SOTA boosting (Nori et al., 2019).
@@
Figure 1: Example of the feature-specific confidence intervals generated by Algorithm 1, compared to the point estimates from EBM (Nori et al., 2019). This visualizes centered feature effects on the UCI Machine Learning Reposity Obesity (Palechor and la Hoz Manotas, 2019) dataset, when predicting weight as the response.
@@
While EBMs provide a clear graphical representation for the effects of each feature, these are still point estimates. The central question we ask for uncertainty quantification can be phrased as this: Given a model design, if we collect a new dataset and retrain the model, how different would our new prediction be? In response to this question, vanilla EBMs (Lou et al., 2012) compute variance estimates via bootstrapping, though this is computationally intensive and only heuristically justified.
@@
A canonical alternative to bootstrapping is to construct confidence and prediction intervals using the asymptotic properties of the model. This is the case with GAMs fitted via the backfitting (but not boosting) algorithm, which enjoy a rich asymptotic theory that enables the construction of confidence and prediction intervals for them (Wahba, 1983).
@@
Unfortunately, the asymptotic behavior of most gradient boosting algorithms (of which EBMs are one), is not known. (Ustimenko et al., 2023) show that a randomized version of boosting converges to a Gaussian process. On the other hand, (Zhou and Hooker, 2019) show that taking an average instead of a sum of trees (in a process which they call Boulevard regularization) converges to a kernel ridge regression, yielding asymptotic normality. This was recently leveraged by (Fang et al., 2025) to construct confidence intervals for the underlying nonparametric function $\mathbf{f}$ and prediction intervals for the labels, with extensions to random dropout and parallel boosting.
@@
Unfortunately, the underlying nonparametric function itself can be highly complicated and uninterpretable in general. Restricting this to the class of GAMs, as with EBMs, yields interpretability. However, the existing Boulevard theory (Zhou and Hooker, 2019) does not directly apply to the class of algorithms we consider here. These results need to be extended to allow different structure matrices (resulting in different kernels) for each feature, along with enforcing the GAM identifiability conditions. While we sketch an algorithm that can be analyzed largely within the framework of (Fang et al., 2025), further algorithmic variants that either employ backfitting directly or employ parallelization methods result in different algorithmic limits requiring separate analysis. We make the following contributions:
  • Asymptotic Theory. We establish central limit theorems for three variations of EBMs (parallelized, backfitting-like parallelized, and sequential), showing that the Boulevard-regularized ensemble converges to a feature-wise kernel ridge regression. The resulting predictions are asymptotically normal and achieve the minimax-optimal MSE of $O(pn^{-2/3})$ for fitting Lipschitz $p$-dimensional GAMs (Yuan and Zhou, 2015). This avoids the curse of dimensionality, significantly improving upon the rates of (Zhou and Hooker, 2019; Fang et al., 2025).
  • Frequentist Inference. By leveraging our proposed CLT for baseline vector $\boldsymbol{\beta}$ and feature-wise predictions $\widehat{\mathbf{f}}^{(k)}$, we extend the heuristic bootstrap to a set of inference toolkits. We provide confidence intervals for each underlying function $\mathbf{f}^{(k)}$ (see Figure 1 for an example), and prediction intervals for the response. This also allows a direct test of variable importance following the methods of (Mentch and Hooker, 2015). We round out those propositions by numerical experiments to back up the validity of our algorithms.
  • Computational Efficiency. The derived procedures above rely on a kernel ridge regression, with a prohibitively expensive $O(n^{3})$ runtime. (Fang et al., 2025) employ the Nyström approximation to produce intervals in $O(np^{2})$ kernel precomputation and $O(p^{2})$ query time, but the recursive implementation of (Musco and Musco, 2017) can be slow in practice. Given that EBMs utilize histogram trees, we instead work out the computations in histogram bin-space. This bin-based implementation runs in an additional lower-order $O(pm^{2})$ time per training round, $O(pm^{3})$ kernel precomputation time, and $O(pm^{2})$ interval query time, where $p$ is the number of features and $m$ the maximum number of bins per-feature. As the number of bins is usually capped at $255$ or $511$, this is independent of the number of samples nn!
@@ 2 Setup and Notation
(Zhou and Hooker, 2019) replace the usual additive update with a moving average. Instead of adding the next tree: $\widehat{\mathbf{f}}_{b}\leftarrow\mathbf{f}_{b-1}+\lambda\mathbf{t}_{b}$, they average over all previous trees: $\widehat{\mathbf{f}}_{b}\leftarrow\frac{b-1}{b}\mathbf{f}_{b-1}+\frac{\lambda}{b}\mathbf{t}_{b}$. This idea of shrinking the previous ensemble as also been adopted by (Ustimenko and Prokhorenkova, 2022; Ustimenko et al., 2023). This ensures convergence to a kernel ridge regression in the limit of infinite boosting rounds, and that the resulting predictions are asymptotically normal. [^2] A particular advantage of Boulevard regularization is that it is robust to overfitting (Fang et al., 2025). Figure 5 shows this behavior for Algorithm 1. EBM achieves this here via tiny learning rates when cycling through features in sequence. We do this while fitting to features in parallel.
@@ 3 Algorithms
1:Input: $\mathbf{X}\in\mathbb{R}^{n\times p}$, $\mathbf{y}\in\mathbb{R}^{n}$, subsample rate $\xi$, learning rate $\lambda$, rounds $T$, truncation level $M>0$.
@@
Averaging yields a kernel ridge regression limit and a CLT, but (as in Boulevard) the fitted signal is asymptotically shrunk, requiring rescaling by $\frac{1+\lambda}{\lambda}$. In Section D, we show that the prediction made by Algorithm 1 converge to a kernel-ridge-regression. Informally:
@@
Though in practice this is not an issue, we explore alternatives below for the sake of completeness. One is to abandon parallelism, and randomly choose a single feature to fit a tree to during each boosting round. This is presented in Algorithm 3 (see Appendix B), which converges to the same fixed point as Algorithm 1, but without the computational gains.
@@
One problem remains. Algorithms 1 and 3 converges to $\frac{\lambda}{1+\lambda}\mathbf{f}\leq\mathbf{f}/2$, and so require rescaling by $\frac{1+\lambda}{\lambda}$. We provide an alternative inspired by the backfitting-like procedure of (Fang et al., 2025) in Algorithm 2 that does not require rescaling.
@@
The original gradient boosting algorithm with Boulevard regularization of (Zhou and Hooker, 2019) converges to a kernel ridge regression of the form $\mathbb{E}[\mathbf{s}(\mathbf{x})]^{\top}(\lambda^{-1}\mathbf{I}+\mathbb{E}[\mathbf{S}])^{-1}\mathbb{E}[\mathbf{S}]\mathbf{y}$. This is similar to our result for Algorithms 1 and 3, but we achieve convergence for each centered univariate function while they achieve convergence only for the overall function $\mathbf{f}$. As a consequence, their method has an MSE of $O(n^{-\frac{1}{d+1}})$ that grows exponentially in dimension. This is not even minimax-optimal for the nonparametric Lipschitz regression setting they are in.
The variants of (Fang et al., 2025) incorporating dropout and parallel training converge to slightly different fixed points, but ultimately achieve the same asymptotic MSE rates exponential in dimension.
@@ 4 Theory
This suggests that ${s}_{n,j}(\mathbf{x})$, the frequency at which training point $\mathbf{x}_{j}$ and test point $\mathbf{x}$ share a leaf, defines a similarity measure. This observation is insightful. In fact, the frequency at which training points $\mathbf{x}_{i},\mathbf{x}_{j}$ share a leaf, $s_{n,j}(\mathbf{x}_{i})$, yields a kernel! [^4] Theorem 3 of (Zhou and Hooker, 2019) shows that $\mathbb{E}\mathbf{S}_{n}$ is p.s.d. with spectral norm no more than 1.
@@
Following (Zhou and Hooker, 2019; Fang et al., 2025), we require that the tree structures be separated from the leaf values, and stabilize after a burn-in period:
@@
The former can be satisfied by refitting leaf values on a hold-out dataset (Wager and Athey, 2017). Non-adaptivity may occur naturally, but can also be enforced by sampling from previous tree structures after a burn-in period. In practice, our algorithms tend to perform well even in the absence of these assumptions.
@@
We control the geometry and size of leaves to ensure locality and stable averages. These assumptions control the MSE of the final estimate – (Zhou and Hooker, 2019) and (Fang et al., 2025) make spiritually similar assumptions, but with different scalings in dimension that do not avoid the curse of dimensionality. Again, we avoid this problem due to the GAM structure of our problem setup:
@@
Unlike (Zhou and Hooker, 2019) and (Fang et al., 2025), the assumptions above are dimension-free. They enforce asymptotic locality of the partition, allowing for control over bias and variance. Assumption 4.5 ensures stability of the structure matrix spectrum, allowing us to recover the minimax-optimal rate later.
@@
As mentioned earlier, the theory for Algorithm 3 is the easiest to prove. With the following observation, this amounts to a straightforward extension of (Zhou and Hooker, 2019). The learner updates a feature’s kernel at random each round, so the tree kernel chosen by Algorithm 3 amounts to $\mathbb{E}\mathbf{S}=\frac{1}{p}\sum_{k=1}^{p}\mathbb{E}[\mathbf{S}^{(k)}]$.
@@
This is a technical assumption novel to the literature that we only use for Algorithm 1. The projector–like condition essentially says the (centered) expected tree kernel for feature $k$ behaves almost recovers only the part of the signal that depends only on $\mathbf{x}^{(k)}$. This is exactly what EBM histogram trees do in the large–$n$ limit: they average locally along $\mathbf{x}^{(k)}$, as leaves get small, Boulevard averaging de–adapts the structures, and centering removes the intercept—so $\mathbf{J}_{n}\mathbf{K}^{(k)}\mathbf{J}_{n}$ nearly projects onto mean–zero functions of $\mathbf{x}^{(k)}$. The conditional orthogonality condition says those per–feature pieces don’t overlap: a centered function of $\mathbf{x}^{(a)}$ is (approximately) orthogonal to a centered function of $\mathbf{x}^{(k)}$ when $a\neq k$ – true under independence and still a good approximation with weak dependence and unaligned, shallow partitions.
@@
One may argue that Assumption 4.7 is unrealistic. If one is not willing to use $\lambda=1/p$ then, Algorithm 2 poses a remedy. This does not rely on Assumption 4.7, and only depends on a weaker assumption below.
@@
Assumption 4.9 is a weaker version of Assumption 4.7. It prevents different features’ kernel spaces from (nearly) coinciding, which would break identifiability/backfitting. This milder assumption buys us a similar finite sample convergence below.
@@
We are now in a position to provide central limit theorems for all three algorithms. Unlike (Zhou and Hooker, 2019) and (Fang et al., 2025), the GAM structure of our problem allows us to relax the rates within the assumptions on tree geometries. While the following lemma allows us to check the Lindeberg-Feller conditions, it also allows us to recover the optimal minimax rate $n^{-2/3}$ for univariate non-parametric Lipschitz regression (Yuan and Zhou, 2015). Below, we use the subscript $E=\{A,B\}$ to label relevant variables in Algorithm 1 and 2.
@@
Normality of the intercept term $\hat{\boldsymbol{\beta}}$ is also established and discussed in Lemma E.15. Similar to (Fang et al., 2025), our CLTs yields a subsequent risk bound:
@@
This result shows that our algorithms achieve the minimax-optimal MSE of $O(pn^{-2/3})$ for fitting Lipschitz $p$-dimensional GAMs (Yuan and Zhou, 2015). This avoids the curse of dimensionality, significantly improving upon the rates of (Zhou and Hooker, 2019; Fang et al., 2025).

### v2

@@ Abstract
Explainable boosting machines (EBMs) are popular “glass-box” models that learn a set of univariate functions using boosting trees. These achieve explainability through visualizations of each feature’s effect. However, unlike linear model coefficients, uncertainty quantification for the learned univariate functions requires computationally intensive bootstrapping, making it hard to know which features truly matter. We provide an alternative using recent advances in statistical inference for gradient boosting, deriving methods for statistical inference as well as end-to-end theoretical guarantees. Using a moving average instead of a sum of trees (Boulevard regularization) allows the boosting process to converge to a feature-wise kernel ridge regression. This produces asymptotically normal predictions that achieve the minimax-optimal MSE for fitting Lipschitz GAMs with $p$ features of $O(p\,n^{-2/3})$, successfully avoiding the curse of dimensionality. We then construct prediction intervals for the response and confidence intervals for each learned univariate function with a runtime independent of the number of datapoints, enabling further explainability within EBMs. Code is available at github.com/hetankevin/ebm-inference.
@@ 1 Introduction
Modern machine learning methods have proven to be highly successful, achieving impressively high performance even for complicated prediction tasks. Still, this comes at a cost. These methods have been critiqued for their lack of transparency, often described as “black-box” models. As such, much work has been done in an attempt to provide explanations for their behavior. [^1] Popular methods include LIME Ribeiro et al. (2016), SHAP Lundberg and Lee (2017), ALE Apley and Zhu (2020) and partial dependence Friedman (2001) plots.Another line of work attempts to convince practitioners to instead use inherently transparent models that are algebraically simple enough to be inspected by humans Rudin (2019). Neither is optimal. Explaining black-box models is a difficult problem Hooker et al. (2021); Fryer et al. (2021), and inherently interpretable models can sacrifice performance.
@@
In light of this, Explainable Boosting Machines (EBM) Lou et al. (2012); Nori et al. (2019) have emerged as a popular “glass-box” model to strike a balance between both. An EBM learns a Generalized Additive Model (GAM) Hastie and Tibshirani (1986): $\mathbf{f}(\mathbf{x})=\sum_{k=1}^{p}\mathbf{f}^{(k)}(\mathbf{x}^{(k)})$, where each $\mathbf{f}^{(k)}$ is assumed to be a univariate function of the $k$-th feature. However, instead of fitting splines, it instead fits a set of univariate gradient boosting regression trees. The tree-based boosting procedure attempts to recover the expressiveness of state-of-the-art algorithms Chen and Guestrin (2016); Ke et al. (2017), while the decomposition into additive components allows the measurement of the marginal contribution of each feature like linear models. While EBMs impose strong structural constraints on the resulting prediction, their ability to reproduce nonlinearities, along with appropriate choices of additive structure, provides comparable performance to SOTA boosting Nori et al. (2019).
@@
Figure 1: Example of the feature-specific confidence intervals generated by Algorithm 1, compared to the point estimates from EBM Nori et al. (2019). This visualizes centered feature effects on the UCI Machine Learning Reposity Obesity Palechor and la Hoz Manotas (2019) dataset, when predicting weight as the response.
@@
While EBMs provide a clear graphical representation for the effects of each feature, these are still point estimates. The central question we ask for uncertainty quantification can be phrased as this: Given a model design, if we collect a new dataset and retrain the model, how different would our new prediction be? In response to this question, vanilla EBMs Lou et al. (2012) compute variance estimates via bootstrapping, though this is computationally intensive and only heuristically justified. Our primary objective is hence well-calibrated frequentist uncertainty quantification for EBMs while retaining their interpretability and scalability.
@@
A canonical alternative to bootstrapping is to construct confidence and prediction intervals using the asymptotic properties of the model. This is the case with GAMs fitted via the backfitting (but not boosting) algorithm, which enjoy a rich asymptotic theory that enables the construction of confidence and prediction intervals for them Wahba (1983). Another widely used alternative for prediction intervals is conformal prediction, which provides finite-sample marginal coverage with minimal assumptions. However, conformal methods do not directly yield feature-wise confidence intervals for the GAM components, which we aim to explore in this work.
@@
Unfortunately, the asymptotic behavior of most gradient boosting algorithms (of which EBMs are one), is not known. Ustimenko et al. (2023) show that a randomized version of boosting converges to a Gaussian process. On the other hand, Zhou and Hooker (2019) show that taking an average instead of a sum of trees (in a process which they call Boulevard regularization) converges to a kernel ridge regression, yielding asymptotic normality. This was recently leveraged by Fang et al. (2025) to construct confidence intervals for the underlying nonparametric function $\mathbf{f}$ and prediction intervals for the labels, with extensions to random dropout and parallel boosting.
@@
Unfortunately, the underlying nonparametric function itself can be highly complicated and uninterpretable in general. Restricting this to the class of GAMs, as with EBMs, yields interpretability. However, the existing Boulevard theory Zhou and Hooker (2019) does not directly apply to the class of algorithms we consider here. These results need to be extended to allow different structure matrices (resulting in different kernels) for each feature, along with enforcing the GAM identifiability conditions. While we sketch an algorithm that can be analyzed largely within the framework of Fang et al. (2025), further algorithmic variants that either employ backfitting directly or employ parallelization methods result in different algorithmic limits requiring separate analysis. Unlike prior Boulevard analyses, EBMs induce feature-specific kernels. This makes the theory substantially more delicate: one must control their interaction under centering and GAM identifiability, and establish asymptotic normality for each additive component, not just the overall predictor. To handle these additional challenges, we make the following contributions:
  • Asymptotic Theory. We establish central limit theorems for three variations of EBMs (parallelized, backfitting-like parallelized, and sequential), showing that the Boulevard-regularized ensemble converges to a feature-wise kernel ridge regression. The resulting predictions are asymptotically normal and achieve the minimax-optimal MSE of $O(pn^{-2/3})$ for fitting Lipschitz $p$-dimensional GAMs Yuan and Zhou (2015). This avoids the curse of dimensionality, significantly improving upon the rates of Zhou and Hooker (2019); Fang et al. (2025).
  • Frequentist Inference. By leveraging our proposed CLT for baseline vector $\boldsymbol{\beta}$ and feature-wise predictions $\widehat{\mathbf{f}}^{(k)}$, we extend the heuristic bootstrap to a set of inference toolkits. We provide confidence intervals for each underlying function $\mathbf{f}^{(k)}$ (see Figure 1 for an example), and prediction intervals for the response. This also allows a direct test of variable importance following the methods of Mentch and Hooker (2015). We round out those propositions by numerical experiments to back up the validity of our algorithms.
  • Computational Efficiency. The derived procedures above rely on a kernel ridge regression, with a prohibitively expensive $O(n^{3})$ runtime. Fang et al. (2025) employ the Nyström approximation to produce intervals in $O(np^{2})$ kernel precomputation and $O(p^{2})$ query time, but the recursive implementation of Musco and Musco (2017) can be slow in practice. Given that EBMs utilize histogram trees, we instead work out the computations in histogram bin-space. This bin-based implementation runs in an additional lower-order $O(pm^{2})$ time per training round, $O(pm^{3})$ kernel precomputation time, and $O(pm^{2})$ interval query time, where $p$ is the number of features and $m$ the maximum number of bins per-feature. As the number of bins is usually capped at $255$ or $511$, this is independent of the number of samples nn!
@@ 2 Setup and Notation
Zhou and Hooker (2019) replace the usual additive update with a moving average. Instead of adding the next tree: $\widehat{\mathbf{f}}_{b}\leftarrow\mathbf{f}_{b-1}+\lambda\mathbf{t}_{b}$, they average over all previous trees: $\widehat{\mathbf{f}}_{b}\leftarrow\frac{b-1}{b}\mathbf{f}_{b-1}+\frac{\lambda}{b}\mathbf{t}_{b}$. This idea of shrinking the previous ensemble as also been adopted by Ustimenko and Prokhorenkova (2022); Ustimenko et al. (2023). This ensures convergence to a kernel ridge regression in the limit of infinite boosting rounds, and that the resulting predictions are asymptotically normal. [^2] A particular advantage of Boulevard regularization is that it is robust to overfitting Fang et al. (2025). Figure 4 shows this behavior for Algorithm 1. EBM achieves this here via tiny learning rates when cycling through features in sequence. We do this while fitting to features in parallel.
@@ 3 Algorithms
1:Input: $\mathbf{X}\in\mathbb{R}^{n\times p}$, $\mathbf{y}\in\mathbb{R}^{n}$, subsample rate $\xi$, learning rate $\lambda$, rounds $B$, truncation level $M>0$.
@@
Averaging yields a kernel ridge regression limit and a CLT, but (as in Boulevard) the fitted signal is asymptotically shrunk, requiring rescaling by $\frac{1+\lambda}{\lambda}$. In Section E, we show that the prediction made by Algorithm 1 converge to a kernel-ridge-regression. Informally:
@@
Though in practice this is not an issue, we explore alternatives below for the sake of completeness. One is to abandon parallelism, and randomly choose a single feature to fit a tree to during each boosting round. This is presented in Algorithm 3 (see Appendix C), which converges to the same fixed point as Algorithm 1, but without the computational gains.
@@
One problem remains. Algorithms 1 and 3 converges to $\frac{\lambda}{1+\lambda}\mathbf{f}\leq\mathbf{f}/2$, and so require rescaling by $\frac{1+\lambda}{\lambda}$. We provide an alternative inspired by the backfitting-like procedure of Fang et al. (2025) in Algorithm 2 that does not require rescaling.
@@
The original gradient boosting algorithm with Boulevard regularization of Zhou and Hooker (2019) converges to a kernel ridge regression of the form $\mathbb{E}[\mathbf{s}(\mathbf{x})]^{\top}(\lambda^{-1}\mathbf{I}+\mathbb{E}[\mathbf{S}])^{-1}\mathbb{E}[\mathbf{S}]\mathbf{y}$. This is similar to our result for Algorithms 1 and 3, but we achieve convergence for each centered univariate function while they achieve convergence only for the overall function $\mathbf{f}$. As a consequence, their method has an MSE of $O(n^{-\frac{1}{d+1}})$ that grows exponentially in dimension. This is not even minimax-optimal for the nonparametric Lipschitz regression setting they are in.
The variants of Fang et al. (2025) incorporating dropout and parallel training converge to slightly different fixed points, but ultimately achieve the same asymptotic MSE rates exponential in dimension.
@@ 4 Theory
This suggests that ${s}_{n,j}(\mathbf{x})$, the frequency at which training point $\mathbf{x}_{j}$ and test point $\mathbf{x}$ share a leaf, defines a similarity measure. This observation is insightful. In fact, the frequency at which training points $\mathbf{x}_{i},\mathbf{x}_{j}$ share a leaf, $s_{n,j}(\mathbf{x}_{i})$, yields a kernel! [^4] Theorem 3 of Zhou and Hooker (2019) shows that $\mathbb{E}\mathbf{S}_{n}$ is p.s.d. with spectral norm no more than 1.
@@
Following Zhou and Hooker (2019); Fang et al. (2025), we require that the tree structures be separated from the leaf values, and stabilize after a burn-in period:
@@
The former can be satisfied by refitting leaf values on a hold-out dataset Wager and Athey (2017). Non-adaptivity may occur naturally, but can also be enforced by sampling from previous tree structures after a burn-in period. In practice, our algorithms tend to perform well even in the absence of these assumptions.
@@
We control the geometry and size of leaves to ensure locality and stable averages. These assumptions control the MSE of the final estimate – Zhou and Hooker (2019) and Fang et al. (2025) make spiritually similar assumptions, but with different scalings in dimension that do not avoid the curse of dimensionality. Again, we avoid this problem due to the GAM structure of our problem setup:
@@
Assumption 4.5 forces trees to perform only local averaging by keeping leaves well-balanced. Assumption 2.1 allows depths as large as $\Omega(\log n)$, which is natural for balanced binary trees.
@@
This is natural: the leaf count cannot exceed the effective histogram resolution. Enforcing at least $n^{2/3-\nu}$ points per leaf rules out overly fine, high-variance partitions.
@@
This is a technically convenient assumption for us to generalize it in random design case. It bounds the effective complexity of the tree class so we can use uniform LLNs and contraction arguments. The assumption still permits an super-polynomial large class in $n$, which is far looser than practice: vanilla EBMs use very shallow trees (typically depth 1–2) on a fixed histogram grid with at most a few hundred bins per feature, so the effective class is much smaller (roughly polynomial in $n$).
Unlike Zhou and Hooker (2019) and Fang et al. (2025), the assumptions above are dimension-free. They enforce asymptotic locality of the partition, allowing for control over bias and variance. Assumption 4.5 ensures stability of the structure matrix spectrum, allowing us to recover the minimax-optimal rate later.
@@
As mentioned earlier, the theory for Algorithm 3 is the easiest to prove. With the following observation, this amounts to a straightforward extension of Zhou and Hooker (2019). The learner updates a feature’s kernel at random each round, so the tree kernel chosen by Algorithm 3 amounts to $\mathbb{E}\mathbf{S}=\frac{1}{p}\sum_{k=1}^{p}\mathbb{E}[\mathbf{S}^{(k)}]$.
@@
Assumption 4.7 is technically convenient but admittedly strong, especially the conditional orthogonality requirement. We therefore emphasize that it is used only to analyze Algorithm 1. As a fallback, Algorithm 2 removes this requirement and needs only Assumption 4.9. Intuitively, the projector-like condition says the centered expected tree kernel for feature $k$ behaves like the orthogonal projector onto mean-zero functions of $\mathbf{x}^{(k)}$, while conditional orthogonality asserts these per-feature subspaces have negligible overlap across $k$ (similar in spirit to “no concurvity” conditions for uniqueness in additive models, e.g., Hastie and Tibshirani, 1986); this can hold exactly under feature independence and can be a reasonable approximation under weak dependence when partitions are shallow and not aligned across features.
@@
$\widetilde{\mathbf{y}}_{k}^{*}\;=\;\mathbf{J}_{n}\,\mathbf{K}^{(k)}\,[\lambda^{-1}\mathbf{I}+\mathbf{J}_{n}\,\mathbf{K}]^{-1}\mathbf{y},\qquad\mathbf{K}\;:=\;\sum_{k=1}^{p}\mathbf{K}^{(k)},$
@@
As discussed before, one may argue that Assumption 4.7 is unrealistic. If one is not willing to use $\lambda=1/p$ then, Algorithm 2 poses a remedy. This does not rely on Assumption 4.7, and only depends on a weaker assumption below.
@@
Assumption 4.9 is a weaker version of Assumption 4.7: instead of requiring (approximate) orthogonality of feature-wise subspaces, it only rules out the degenerate case where different features induce nearly identical smoothing operators in expectation. We verified this assumption empirically with plausible results. See Table 10. This milder assumption is sufficient for establishing a similar finite-sample convergence result below.
@@
We are now in a position to provide central limit theorems for all three algorithms. Unlike Zhou and Hooker (2019) and Fang et al. (2025), the GAM structure of our problem allows us to relax the rates within the assumptions on tree geometries. While the following lemma allows us to check the Lindeberg-Feller conditions, it also allows us to recover the optimal minimax rate $n^{-2/3}$ for univariate non-parametric Lipschitz regression Yuan and Zhou (2015). Below, we use the subscript $E=\{A,B\}$ to label relevant variables in Algorithm 1 and 2.
@@
Normality of the intercept term $\hat{\boldsymbol{\beta}}$ is also established and discussed in Lemma F.15. Similar to Fang et al. (2025), our CLTs yields a subsequent risk bound:
@@
This result shows that our algorithms achieve the minimax-optimal MSE of $O(pn^{-2/3})$ for fitting Lipschitz $p$-dimensional GAMs Yuan and Zhou (2015). This avoids the curse of dimensionality, significantly improving upon the rates of Zhou and Hooker (2019); Fang et al. (2025).